OpenStudy (happykiddo):

How do I figure out the number of atoms in element, when I know the molar mass is 3.30 ???

OpenStudy (anonymous):

whats the element ?

OpenStudy (happykiddo):

Krypton

OpenStudy (anonymous):

so you have 3.3 moles of Krypton, and you want to find how many atoms are there, is tha t right ?

OpenStudy (happykiddo):

Yes

OpenStudy (anonymous):

replace mole wid Avogadro number

OpenStudy (happykiddo):

Ok

OpenStudy (anonymous):

1 mole = 6.022 × 10^23 items

OpenStudy (anonymous):

just like, 1 dozen = 12 items

OpenStudy (anonymous):

3.3 moles of Kr = 3.3 x 6.022 × 1023 atoms

OpenStudy (happykiddo):

Thanks, a bunch. :)

OpenStudy (happykiddo):

We're did you get1023

OpenStudy (happykiddo):

Please answer me , I need some clarification .

OpenStudy (anonymous):

3.3 moles of Kr = \(3.3 \times 6.022 × 10^{23}\) atoms

OpenStudy (anonymous):

its like that

OpenStudy (anonymous):

\(1\ mole = 6.022 \times 10^{23} \) atoms

OpenStudy (happykiddo):

Oh oops ok thank you very much. :)

OpenStudy (anonymous):

wlc :)

OpenStudy (happykiddo):

So the answer is 19x10 to 23 atom According to significant figures

OpenStudy (anonymous):

you wanto show it in scientific notation ?

OpenStudy (happykiddo):

So the answer is 19 x 10 to 23 power According to significant figures

OpenStudy (anonymous):

\(3.3 \times 6.022 × 10^{23}\) \(19.8726 \times 6.022 × 10^{23}\) \(1.98726 \times 6.022 × 10^{24}\)

OpenStudy (anonymous):

thats scientific notation, we must change that 19.xxxx to 1.9xxxx

OpenStudy (happykiddo):

Oh yeah forget .....well really once again thank you.

OpenStudy (anonymous):

yw